One-fifth of the sum of two consecutive integers is 1 less than the smaller integer. Find the integers.
Answer by stanbon(75887)

One-fifth of the sum of two consecutive integers is 1 less than the smaller integer. Find the integers.
---------------
Let the consecutive integers be x and x+1.
EQUATION:
(1/5)(x+x+1)+1=x
2x+1=5(x-1)
2x+1=5x-5
3x=6
x=2
x+1=3
----------
Cheers,
Stan H.